Consider a recurrence relation an = 3an-1 for n = 1,2,3,4,... and suppose a=4. Calculate a5.


1
Expert's answer
2021-11-24T13:47:19-0500
"a_5=3a_4=3(3a_3)=9(3a_2)=27(3a_1)=81(4)=324"

We have geometric sequence


"a=a_1=4, q=\dfrac{a_n}{a_{n-1}}=3, n\geq2"

Then

"a_n=a(q)^{n-1}"

"a_5=4(3)^{5-1}=324"